21
Nov
2010
08:01

There is currently a problem with one component of a core switch cluster in Maidenhead. We will shortly be resetting it. If this does not restore normal operation, we will need to reboot the entire cluster, which will entail a minute or so downtime for all services traversing the cluster.

Update 10:20

The switch element in question (beatrix2) has suffered complete hardware failure. Fortunately most of the services it provides are dual-fed; those not have been re-routed. Replacement hardware has been ordered and will hopefully be installed on Tuesday.

Posted by admin

09
Nov
2010
14:06

We have just had a short outage on our BT connections. Connections have now restored. We are awaiting an explanation.

Posted by admin

Current Status

    GoodNo current issues

Status Archive